# Unraveling additional quantum many-body scars of the spin-$1$ XYXY model with Fock-space cages and commutant algebras

### v1

@@ Abstract
Quantum many-body scars (QMBS) represent a mechanism for weak ergodicity breaking, characterized by the coexistence of atypical non-thermal eigenstates within an otherwise thermalizing many-body spectrum. In this work, we revisit the spin-$1$ XYXY model on a periodic chain and construct several new families of exact scar eigenstates embedded within its extensively degenerate manifolds that owe their origins to an interplay of $U(1)$ magnetization conservation and chiral symmetries. We go beyond previously studied towers of states and first identify a novel set of interference-protected eigenstates resembling Fock space cage states, where destructive interference confines the wave function to sparse subgraphs of the Fock space. These states exhibit subextensive entanglement entropy, and when subjected to a transverse magnetic field, form equally spaced ladders whose coherent superpositions display long-lived fidelity oscillations. We further reveal a simpler organizing principle behind these nonthermal states by using the commutant algebra framework, in particular by showing that they are simultaneous eigenstates of non-commuting local operators. Moreover, in doing so, we uncover two more novel families of exact scars: a tower of volume-entangled states, and a set of mirror-dimer states with some free local degrees of freedom. Our results illustrate the power and interplay of interference-based and algebraic mechanisms of non-ergodicity, offering systematic routes to identifying and classifying QMBS in generic many-body quantum systems.
@@ I Introduction
Beyond these strong-ETH-violating paradigms, recent years have seen the emergence of a novel class of systems that break thermalization in a much weaker sense. These systems, which are otherwise chaotic and nonintegrable, host a small subset of non-thermal eigenstates embedded within a thermal spectrum, a phenomenon now known as quantum many-body scars (QMBS) [15, 16, 17, 18, 19]. Scar states often occur in quasiparticle towers of states [20, 19], which exhibit anomalously low entanglement entropy (EE) compared to typical eigenstates at similar energies, although there are many other kinds of scars, including volume-law entangled ones [21, 22, 23, 24, 25, 26] which are nevertheless atypical in different ways. The existence of these states in the spectrum can lead to non-thermal dynamics in the expectation values of local observables when the system is quenched from an initial state that has a high overlap with the scar states. QMBS have been experimentally observed in the PXP model [27, 28], both exact [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49] and approximate QMBS [50, 16, 51, 52, 53, 54, 55, 56, 57, 58, 59] have also been identified in a wide range of models.
@@
Among the burgeoning landscape of exact scar-hosting systems, those featuring extensive zero-energy modes [50, 60, 61, 62, 63, 64, 65, 66] have emerged as particularly intriguing. In such systems, symmetries-most commonly chiral and/or spectral-reflection symmetries-enforce a macroscopically large manifold of many-body eigenstates pinned at exactly zero energy [60]. While these zero-energy subspaces, typically located at the center of the spectrum, might be expected to host thermal states, several works [67, 61, 36, 62, 63, 68, 69, 25, 70, 22] have demonstrated that they can also harbor non-thermal, scar-like states that arise as specific linear combinations of the zero-modes. These scattered examples of scars within models possessing extensive zero-modes hint at a rich underlying structure. Yet, these are hard to construct precisely because of the extensive degeneracy of the manifold of states in which they lie, which leads conventional methods of scar identification, such as anomalous entanglement entropy, to fail. Systematic and comprehensive construction mechanisms to detect such scars are mostly lacking. Moreover, the structures of these manifolds are themselves largely unexplored in models with multiple symmetries such as $U(1)$ [71] along with chiral or inversion symmetries.
@@ II.1 Model
Here, $S^{\alpha}_{i}(\alpha{=}x,y,z)$ are spin-$1$ operators acting on local basis states $|m\rangle_{i}{\in}\{|{-}1\rangle,|0\rangle,|1\rangle\}$ at site $i$ satisfying $S^{z}|m\rangle{=}m|m\rangle$. The Hamiltonian $H_{0}$ consists of two parts: a nearest-neighbor $XY$ exchange interaction $H_{XY}$, the strength of which is controlled by the coupling constant $J$, and a Zeeman term $H_{z}$ corresponding to a uniform magnetic field of strength $h$ along the $z$-axis. The Hamiltonian $H_{0}$ conserves the total magnetization $S_{\rm tot}^{z}{=}\sum_{i=1}^{2L}S^{z}_{i}$, leading to a global $U(1)$ symmetry that partitions the Hilbert space into magnetization sectors $M{\in}\{{-}2L,{-}2L{+}1,{\cdots},2L{-}1,2L\}$. Additionally, the model has spin-inversion, reflection, and translational symmetries.
@@
Interestingly, the exchange term $H_{XY}$ admits an extensive number of exact zero-energy eigenstates as evident from its density of states shown in Fig. 1, which exhibits a sharp peak at E=0E{=}0. This high degeneracy forms the basis of our construction of analytically tractable exact eigenstates. Our goal is to quantify this degeneracy and explicitly construct several distinct families of exact eigenstates within the zero-energy manifold of $H_{XY}$. Throughout this work, when we use the phrase “zero-energy eigenstates,” we mean the eigenstates have zero energy with respect to $H_{XY}$. When the Zeeman field $H_{z}$ is turned on, these states acquire finite energies under $H_{0}$ that are integer multiples of the field $h$ and, as we will show, display characteristic features of QMBS. Even in the presence of the Zeeman field $H_{z}$, the degeneracy of states, now at finite energies, remains extensive.
@@
Before we proceed further, we note an important and subtle point regarding the integrability of $H_{0}$. Previous studies [83, 32, 33] have established that $H_{0}$ exhibits Poissonian level-statistics in even magnetization sectors, indicating a symmetry, which has been identified to be a hidden non-local $SU(2)$ symmetry, which is additionally believed to lead to integrability of those sectors [33]. This structure complicates the direct interpretation of the eigenstates we construct as true QMBS, since QMBS are typically defined in non-integrable systems. Nevertheless, we show that our interpretation of QMBS as simultaneous eigenstates of non-commuting operators naturally evades this problem, since this allows us to construct many kinds of perturbations $V_{\rm pert}$ that break the non-local symmetry and remove integrability while still preserving our constructed eigenstates as exact scar eigenstates of the perturbed Hamiltonian $H_{\rm scar}{=}H_{0}{+}V_{\rm pert}$. The particular forms of $V_{\rm pert}$ vary depending on the scar family and will be specified in the relevant sections.
@@
Readers might also notice that, unlike previous studies of the spin-$1$ XYXY model in the QMBS literature [32], we do not include the single-ion anisotropy term $\sum_{j}{(S_{j}^{z})^{2}}$ or the next-nearest-neighbor $XY$ exchange term $J_{3}\sum_{i=1}^{2L}(S^{x}_{i}S^{x}_{i+3}+S^{y}_{i}S^{y}_{i+3})$ in the definition of the model. This choice is deliberate, as such perturbations were included as the appropriate $V_{\rm pert}$ for stabilizing the set of previously-obtained QMBS while breaking the fine-tuned symmetries of $H_{0}$. As we will discuss, the exact eigenstate families we construct respond differently to those perturbations, and hence are not always the appropriate $V_{\rm pert}$ for the QMBS we obtain. In particular, the tower of states of Sec. III.2 remain exact eigenstates upon adding the anisotropy term, but are destroyed by the next-nearest-neighbor $XY$ exchange, whereas the volume-entangled and mirror-dimer families of Sec. IV exhibit the opposite behavior. We defer the detailed discussion of these properties to the corresponding sections; here we simply note that the Hamiltonian in Eq. (1) provides a common baseline for all subsequent constructions.
@@ II.2 Chiral symmetry and extensive zero modes
In addition to the total magnetization, $H_{XY}$ also possesses a chiral symmetry that plays a central role in generating its extensive zero-energy degeneracy. For the chains of even length $2L$ under consideration, this symmetry is generated by the unitary and Hermitian operator
@@
where $\iota{=}\sqrt{{-}1}$, that satisfies the anticommutation relation $\{H_{XY},\mathcal{\hat{C}}\}{=}0$. This implies the energy spectrum of $H_{XY}$ is symmetric around zero, as seen in Fig. 1.
@@
First, we can see that the Fock-space graph is bipartite. To understand how the chiral symmetry shapes this structure, consider the action of the operator $\mathcal{P}_{i}^{z}$ on the local basis states: $\mathcal{P}_{i}^{z}|m\rangle_{i}{=}(-1)^{m{+}1}|m\rangle_{i}$, i.e., it flips the sign of a state $|0\rangle_{i}$ while leaving $|{\pm}1\rangle_{i}$ unchanged. Therefore, the chiral operator $\mathcal{\hat{C}}$ assigns a $\mathbb{Z}_{2}$ value to each basis state based on the parity of the number of $|0\rangle$’s on its even-numbered sites. Crucially, the action of $H_{XY}$ flips the chiral parity of any basis state, since for any computational basis state $|\Lambda\rangle$ satisfying $\mathcal{\hat{C}}|\Lambda\rangle{=}\lambda|\Lambda\rangle$ with $\lambda{=}{\pm}1$, the anticommutation of $\mathcal{\hat{C}}$ and $H_{XY}$ gives $\mathcal{\hat{C}}(H_{XY}|\Lambda\rangle){=}{-}\lambda(H_{XY}|\Lambda\rangle)$. Thus, $H_{XY}$ connects configurations with opposite $\mathcal{\hat{C}}-$parity, implying that the Fock-space graph is bipartite: its nodes can be divided into two disjoint sublattices
@@
Using a generating-function method (see Appendix A for details), we compute these quantities analytically. Ultimately, the Hilbert-space dimension in a given magnetization sector $M$, denoted by $|\mathcal{D}_{2L}^{M}|$, is obtained as the coefficient of $x^{M}$ in the expansion of $(x{+}1{+}x^{-1})^{2L}$, leading to a trinomial coefficient
$|\mathcal{D}_{2L}^{M}|{=}\sum_{p=\max(0,M)}^{\lfloor(M{+}2L)/2\rfloor}\binom{2L}{p}\binom{2L{-}p}{M{+}2L-2p}$. (5)
Similarly, the lower bound on the number of zero modes is given by the coefficient of $x^{M}$ in $(x^{2}{+}1{+}x^{-2})^{L}$, which is non-zero only for even $M$:
@@
This extensive degeneracy at zero energy of $H_{XY}$, which lies precisely at the middle of its spectrum, opens the possibility of constructing non-generic superpositions of the degenerate eigenstates with features reminiscent of QMBS, which we explore below. Multiple recent works have shown that such zero-mode manifolds can host interesting eigenstates, with either area-law [67, 70] or volume-law entanglement [23, 22, 24, 25, 26], demonstrating the possibility of interesting kinds of nonthermal phenomena in the manifold.
@@ II.3 Previously known scar eigenstates in the zero-energy manifold of $H_{XY}$
Another family, known as bond-bimagnon states $|\mathbb{B}_{n}^{\prime}\rangle$, was introduced in Ref. [32] and later examined in detail in Ref. [33]. These states are given by
@@
These examples highlight the rich structure of the zero-energy manifold $H_{XY}$ and motivate a broader search for hidden nonthermal eigenstates within it. In the remainder of this work, we construct several previously unreported families of exact eigenstates, arising from mechanisms of destructive interference in Fock space and from algebraic organization through commutant structures, thereby substantially extending the known landscape of exact scars in the spin-$1$ XYXY model.
@@ III Novel family of Fock-space cage eigenstates
In this section, we introduce a class of exact zero-energy eigenstates of $H_{XY}$ that are confined to compact, dynamically isolated subgraphs within its many-body Fock space. These states, which we refer to as Fock-space cage (FSC) [72, 71, 73, 74] states, arise from local interference that prevents the spreading of wave function amplitude under the action of the $H_{XY}$, leading to perfectly Fock-space localized eigenstates (although not in real-space) in a disorder-free, translation-invariant system. A crucial property that enables such interference-based localization is that $H_{XY}$ connects basis states with equal transition amplitudes, making its Fock-space graph effectively binary, with off-diagonal elements taking only values 0 or $1$. As a result, localization of states in Fock space can arise purely from sign cancellations rather than fine-tuned amplitudes, making them much more likely to occur. While such interference-induced localization has recently been explored in constrained settings [72, 71, 73, 74], our finding reveals that a similar mechanism can emerge in a fully unconstrained model without any fine-tuned geometry. In particular, previously reported exact scar states of the spin-$1$ XYXY model given in Eqs. (8) and (9) also lend themselves to an interpretation in terms of FSCs as illustrated in Appendix B.1. Moreover, as we demonstrate and focus on below, FSCs are effective in uncovering new scar families in the spin-$1$ XYXY chain.

### v2

# Additional quantum many-body scars of the spin-$1$ XYXY model with Fock-space cages and commutant algebras

@@ Abstract
Quantum many-body scars (QMBS) represent a mechanism for weak ergodicity breaking, characterized by the coexistence of atypical non-thermal eigenstates within an otherwise thermalizing many-body spectrum. In this work, we revisit the spin-$1$ XYXY model on a periodic chain and construct several new families of exact scar eigenstates embedded within its extensively degenerate manifolds that owe their origins to an interplay of $U(1)$ magnetization conservation and chiral symmetries. We go beyond previously studied towers of states and first identify a novel set of interference-protected eigenstates resembling Fock space cage states, where destructive interference confines the wave function to sparse subgraphs of the Fock space. These states exhibit subextensive entanglement entropy, and when subjected to a transverse magnetic field, form equally spaced states whose coherent superpositions display long-lived fidelity oscillations. We further reveal a simpler organizing principle behind these nonthermal states by utilizing the commutant algebra framework, specifically by demonstrating that they are simultaneous eigenstates of non-commuting local operators. Moreover, in doing so, we uncover two more novel families of exact scars: a tower of volume-entangled states, and a set of mirror-dimer states with some free local degrees of freedom. Our results illustrate the power and interplay of interference-based and algebraic mechanisms of non-ergodicity, offering systematic routes to identifying and classifying QMBS in generic many-body quantum systems.
@@ I Introduction
Beyond these strong-ETH-violating paradigms, recent years have seen the emergence of a novel class of systems that break thermalization in a much weaker sense. These systems, which are otherwise chaotic and nonintegrable, host a small subset of non-thermal eigenstates embedded within a thermal spectrum, a phenomenon now known as quantum many-body scarring [15, 16, 17, 18, 19]. Quantum many-body scars (QMBS) states often occur in quasiparticle towers of states [17, 19], which exhibit anomalously low entanglement entropy (EE) compared to typical eigenstates at similar energies. However, there are many other kinds of scars, including volume-law entangled ones [20, 21, 22, 23, 24, 25] which are nevertheless atypical in different ways. The existence of these states in the spectrum can lead to non-thermal dynamics in the expectation values of local observables when the system is quenched from an initial state that has a high overlap with the scar states. QMBS have been experimentally observed in Rydberg atom arrays that approximately simulate the PXP model [26, 27]. Following this, both exact [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48] and approximate QMBS [16, 49, 50, 51, 52, 53, 54, 55, 56, 57, 15] have been identified in a wide range of models.
@@
Among the burgeoning landscape of exact scar-hosting systems, those featuring extensive zero-energy modes [15, 58, 59, 60, 61, 62, 63, 64] have emerged as particularly intriguing. In such systems, symmetries-most commonly chiral and/or spectral-reflection symmetries-enforce a macroscopically large manifold of many-body eigenstates to be pinned at exactly zero energy [58]. While these zero-energy subspaces, typically located at the center of the spectrum, might be expected to host thermal states, several works [65, 59, 35, 60, 61, 37, 66, 24, 67, 21] have demonstrated that they can also harbor non-thermal, scar-like states that arise as specific linear combinations of the zero-modes. These scattered examples of scars within models possessing extensive zero-modes hint at a rich underlying structure. Yet, these are hard to construct precisely because of the extensive degeneracy of the manifold of states in which they lie, which leads conventional methods of scar identification, such as anomalous entanglement entropy, to fail. We note that Ref. [60] proposed an algorithm to deal with this degeneracy and identify scars by finding the lowest entangled states in this manifold, which successfully identifies many interesting states such as the exact MPS states in the zero-energy manifold of the PXP model [65, 67]. In fact, they also go on to conjecture that every local Hamiltonian with such a zero-energy manifold has an area-law entangled state within that manifold, although it is not guaranteed to be an MPS with a finite bond dimension. While such methods are certainly useful in finding exact and approximate area-law scars in this manifold, as mentioned above and as we shall show below, there exist scars in this manifold that have higher entanglement but yet fairly simple structures, such as quasiparticle towers or even simple volume-law entangled states, and it is not clear if such algorithms can be generalized to capture these states. This gap highlights the need for systematic and analytic construction methods capable of uncovering more kinds of nonthermal states or subspaces. Moreover, the structures of these manifolds are themselves largely unexplored in models with multiple symmetries such as $U(1)$ [68] along with chiral or inversion symmetries, and there might be additional surprises hiding there.
@@ II.1 Model
Here, $S^{\alpha}_{i}(\alpha{=}x,y,z)$ are spin-$1$ operators acting on local basis states $|m\rangle_{i}{\in}\{|{-}1\rangle,|0\rangle,|1\rangle\}$ at site $i$ satisfying $S^{z}|m\rangle{=}m|m\rangle$. The Hamiltonian $H_{0}$ consists of two parts: a nearest-neighbor $XY$ exchange interaction $H_{XY}$, the strength of which is controlled by the coupling constant $J$, and a Zeeman term $H_{z}$ corresponding to a uniform magnetic field of strength $h$ along the $z$-axis. The Hamiltonian $H_{0}$ conserves the total magnetization $S_{\rm tot}^{z}{=}\sum_{i=1}^{2L}S^{z}_{i}$, leading to a global $U(1)$ symmetry that partitions the Hilbert space into magnetization sectors $M{\in}\{{-}2L,{-}2L{+}1,{\cdots},2L{-}1,2L\}$. Additionally, the model has reflection, and translational symmetries.
@@
Interestingly, the exchange term $H_{XY}$ admits an extensive number of exact zero-energy eigenstates as evident from its density of states shown in Fig. 1, which exhibits a sharp peak at E=0E{=}0. This high degeneracy forms the basis of our construction of analytically tractable exact eigenstates. Our goal is to quantify this degeneracy and explicitly construct several distinct families of exact eigenstates within the zero-energy manifold of $H_{XY}$. Throughout this work, when we use the phrase “zero-energy eigenstates,” we mean the eigenstates have zero energy with respect to $H_{XY}$. When the Zeeman field $H_{z}$ is turned on, these states acquire finite energies under $H_{0}$ equal to $hM$ that are integer multiples of the field $h$ and, as we will show, display characteristic features of QMBS. Even in the presence of the Zeeman field $H_{z}$, the degeneracy of states, now at finite energies, remains extensive.
@@
Before we proceed further, we note an important and subtle point regarding the integrability of $H_{0}$. Previous studies [80, 31, 32] established that $H_{0}$ exhibits Poissonian level-statistics in even magnetization sectors, indicating a symmetry, which has been identified to be a hidden non-local $SU(2)$ symmetry, which was additionally believed to lead to integrability of those sectors [32]. However, once this non-local $SU(2)$ symmetry is resolved, the level statistics match those of the Gaussian Orthogonal Ensemble (GOE) of the random-matrix theory, demonstrating that the spin-$1$ XYXY model is not integrable [81]. Since QMBS are typically defined only in non-integrable systems after resolving all symmetries, the eigenstates we construct are true QMBS of the spin-$1$ XYXY model, even without the addition of any other perturbations. Moreover, we show that our interpretation of QMBS as simultaneous eigenstates of non-commuting operators allows us to construct many kinds of perturbations $V_{\rm pert}$ that readily break the non-local symmetry while still preserving our constructed eigenstates as exact scar eigenstates of the perturbed Hamiltonian $H_{\rm scar}{=}H_{0}{+}V_{\rm pert}$, which is even more clearly non-integrable. The particular forms of $V_{\rm pert}$ vary depending on the scar family and will be specified in the relevant sections.
@@
Readers might also notice that, unlike previous studies of the spin-$1$ XYXY model in the QMBS literature [31], we omit the single-ion anisotropy term $\sum_{i=1}^{2L}{(S_{i}^{z})^{2}}$ or the next-to-next-nearest-neighbor $XY$ exchange term $J_{3}\sum_{i=1}^{2L}(S^{x}_{i}S^{x}_{i+3}{+}S^{y}_{i}S^{y}_{i+3})$ in the definition of the model. This choice is deliberate, as such perturbations were included as the appropriate $V_{\rm pert}$ for stabilizing the set of previously-obtained QMBS while breaking the fine-tuned symmetries of $H_{0}$ (Note that the analogous next-nearest-neighbor $XY$ exchange term $J_{2}\sum_{i=1}^{2L}(S^{x}_{i}S^{x}_{i+2}{+}S^{y}_{i}S^{y}_{i+2})$ does not preserve either the previously obtained scars [82] or the scars we unravel in this work.). As we will discuss, the exact eigenstate families we construct respond differently to those perturbations, and hence are not always the appropriate $V_{\rm pert}$ for the QMBS we obtain. In particular, the tower of states of Sec. III.2 remain exact eigenstates upon adding the anisotropy term, but are destroyed by the next-nearest-neighbor $XY$ exchange, whereas the volume-entangled and mirror-dimer families of Sec. IV exhibit the opposite behavior. We defer the detailed discussion of these properties to the corresponding sections; here we simply note that the Hamiltonian in Eq. (1) provides a common baseline for all subsequent constructions.
@@ II.2 Chiral symmetry and extensive zero modes
In addition to the total magnetization, $H_{XY}$ also possesses a chiral symmetry that plays a central role in generating its extensive zero-energy degeneracy. For the chains of even length $2L$ under consideration, this symmetry is generated by the unitary and Hermitian operator, whose action is non-trivial only on the even-numbered sites,
@@
where $\iota{=}\sqrt{{-}1}$, that satisfies the anticommutation relation $\{H_{XY},\mathcal{\hat{C}}\}{=}0$. This implies the energy spectrum of $H_{XY}$ is symmetric around zero, i.e., has particle-hole symmetry, as seen in Fig. 1.
@@
First, we can see that the Fock-space graph is bipartite. To understand how the chiral symmetry shapes this structure, consider the action of the operator $\mathcal{P}_{i}^{z}$ on the local basis states: $\mathcal{P}_{i}^{z}|m\rangle_{i}{=}(-1)^{m{+}1}|m\rangle_{i}$, i.e., it flips the sign of the state $|0\rangle_{i}$ while leaving $|{\pm}1\rangle_{i}$ unchanged. Therefore, the chiral operator $\mathcal{\hat{C}}$ assigns a $\mathbb{Z}_{2}$ value to each basis state based on the parity of the number of $|0\rangle$’s on its even-numbered sites. Crucially, the action of $H_{XY}$ flips the chiral parity of any basis state, since for any computational basis state $|\Lambda\rangle$ satisfying $\mathcal{\hat{C}}|\Lambda\rangle{=}\lambda|\Lambda\rangle$ with $\lambda{=}{\pm}1$, the anticommutation of $\mathcal{\hat{C}}$ and $H_{XY}$ gives $\mathcal{\hat{C}}(H_{XY}|\Lambda\rangle){=}{-}\lambda(H_{XY}|\Lambda\rangle)$. Thus, $H_{XY}$ connects configurations with opposite $\mathcal{\hat{C}}-$parity, implying that the Fock-space graph is bipartite: its nodes can be divided into two disjoint sublattices
@@
Using a generating-function method (see Appendix A for details), we compute these quantities analytically. Ultimately, the Hilbert-space dimension in a given magnetization sector $M$, denoted by $\mathcal{D}_{2L}^{M}$, is obtained as the coefficient of $x^{M}$ in the expansion of $(x{+}1{+}x^{-1})^{2L}$, leading to the trinomial coefficient
$\mathcal{D}_{2L}^{M}{=}\sum_{p=\max(0,M)}^{\lfloor(M{+}2L)/2\rfloor}\binom{2L}{p}\binom{2L{-}p}{M{+}2L-2p}$. (5)
Similarly, the lower bound on the number of zero modes is given by the coefficient of $x^{M}$ in $(x^{2}{+}1{+}x^{-2})^{L}$, which is non-zero only for even $M$, and given by $\mathcal{D}_{L}^{M/2}$, i.e.
@@
This extensive degeneracy at zero energy of $H_{XY}$, which lies precisely at the middle of its spectrum, opens the possibility of constructing non-generic superpositions of the degenerate eigenstates with features reminiscent of QMBS [36], which we explore below. Multiple recent works have shown that such zero-mode manifolds can host interesting eigenstates, with either area-law [65, 67] or volume-law entanglement [22, 21, 23, 24, 25], demonstrating the possibility of interesting kinds of nonthermal phenomena in the manifold.
@@ II.3 Previously known scar eigenstates in the zero-energy manifold of $H_{XY}$
Another family, known as bond-bimagnon states $|\mathbb{B}_{n}^{\prime}\rangle$, was introduced in Ref. [31] and later examined in detail in Ref. [32]. These states, given by
@@
These examples highlight the rich structure of the zero-energy manifold of $H_{XY}$ and motivate a broader search for hidden nonthermal eigenstates within it. In the remainder of this work, we construct several previously unreported families of exact eigenstates, arising from mechanisms of destructive interference in Fock space and from algebraic organization through commutant structures, thereby substantially extending the known landscape of exact scars in the spin-$1$ XYXY model.
@@ III Novel family of Fock-space cage eigenstates
In this section, we introduce a class of exact zero-energy eigenstates of $H_{XY}$ that are confined to compact, dynamically isolated subgraphs within its many-body Fock space. These states, which we refer to as Fock-space cage (FSC) [69, 68, 70, 71] states, arise from local interference that prevents the spreading of wave function amplitude under the action of the $H_{XY}$, leading to perfectly Fock-space localized eigenstates (although not in real-space) in a disorder-free, translation-invariant system. A crucial property that enables such interference-based localization is that $H_{XY}$ connects basis states with equal transition amplitudes, making its Fock-space graph effectively binary, with off-diagonal elements, when appropriately normalized, taking only values 0 or $1$. As a result, localization of states in Fock space can arise purely from sign cancellations rather than fine-tuned amplitudes, increasing their likelihood of occurrence. While such interference-induced localization has recently been explored in constrained settings [69, 68, 70, 71], our finding reveals that a similar mechanism can emerge in a fully unconstrained model without any fine-tuned geometry. In particular, previously reported exact scar states of the spin-$1$ XYXY model given in Eqs. (8) and (9) also lend themselves to an interpretation in terms of FSCs as illustrated in Appendix B.1. Moreover, as we demonstrate and focus on below, FSCs are effective in uncovering new scar families in the spin-$1$ XYXY chain.